ADNAN MUHAMMED

Grade 12,

evaluate ∫ cot^-1( sin 2x / 1 – cos 2x) dx

evaluate
∫  cot^-1( sin 2x / 1 – cos 2x) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int cot^{-1}(\frac{sin2x}{1-cos2x})dx
I = \int cot^{-1}(\frac{2sinx.cosx}{2sin^2x})dx
I = \int cot^{-1}(\frac{cosx}{sinx})dx
I = \int cot^{-1}(cotx)dx
I = \int xdx
I = \frac{x^2}{2}+c